# Large Differences Between Stochastic and Deterministic Kinetics in a Simple Autocatalytic Reaction Network

### v2

@@ I Introduction
The stochastic delay and the fluctuations around the mean are especially significant when the system is initially free of $\mathrm{B}$ and when $k_{1}\ll k_{2}$. For this reason, unlike in Ref. Schuster (2019), we do not use SSA to compute the stochastic dynamics of the system—Gillespie’s standard algorithm becomes ineffective when reaction rates differ greatly, and advanced variants such as tau-leaping are not fully equivalent to the CME. Therefore, we derive an explicit analytical solution of the CME for the reactions (1)–(2).
@@ II Theoretical Framework
We assume the system under consideration to be spatially homogeneous and closed, with no exchange of molecules with the environment (e.g., well-stirred reagents in a batch reactor). Consequently, the total molecule number is conserved:
@@ II.1 Irreversible reactions
Both the Finke–Watzky model (1)–(2) and the two other sets of irreversible reactions considered here are one-step processes. Their Chemical Master Equation (CME) takes the general form van Kampen (2007):
@@
for $n=1,\dots,N-1$. $P_{0}(t)$ follows from normalization or directly from Eq. (6) at $n=0$. Here $C_{NN}=1$. Finally, since $r_{0}=0$ and $r_{n}>0$ for $n\geq 1$, the steady-state solution is
@@
The general analytical solution (10)–(11) of Eqs. (6) has appeared in various contexts and has been derived independently by different authors Allnatt and Jacobs (1968); Arslan and Laurenzi (2008); Lee and Kim (2012); Rudnicki (2014); Jedrak (2014). In contrast, by "explicit analytical solution" we mean here not the general form (10)–(11) of the solution, but the most concise analytical form of both $P_{n}(t)$ and its lowest moments that can be obtained for a specific choice of the functional form of the coefficients $r_{n}$.
@@ II.2 Reversible reactions
For reversible bimolecular reactions in general, and for (3)–(4) in particular, the functional form of $\mathbf{P}(t)$ can no longer be determined analytically, and we must resort to numerical methods. This may involve either SSA simulations or a direct, brute-force, exact numerical solution of the system of ODEs (14) or a much more sophisticated method of Ref. Smith and Shahrezaei (2015).
@@ III Results
In this section, we present our results. First, we derive the explicit analytical solution of the CME for the original formulation of the Finke–Watzky model (FWM) (1)–(2). We then use this solution to analyze the differences between stochastic and deterministic kinetics for selected choices of model parameters and initial conditions (Subsection III.1.2). The approximate approach based on a truncated set of moment equations is discussed in Subsection III.1.4, whereas in Subsection III.1.5 we analyze the origin of the high sensitivity of the CME solution to the choice of the initial state.
@@
$\tilde{M}\equiv M+\nu,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \nu\equiv\frac{k_{1}}{k_{2}},$ (20)
@@ III.1.1 Probability distribution and its two lowest moments
Analytical expressions for the moments (III.1.1)–(III.1.1) reduce the double sum (which arises when calculating the moment directly from its definition using $P_{n}(t)$ given by Eq. 22) to a single sum, lowering the computational cost. A caveat is that numerical evaluation of Eqs. (III.1.1) and (III.1.1) requires arbitrary-precision arithmetic: for $M=100$ the absolute values of the coefficients $E^{(1)}_{l}$ and $E^{(2)}_{l}$ reach $10^{30}$ and $10^{33}$, respectively, so small relative round-off errors at short times can cause large absolute errors in the moments.
@@ III.1.2 When is the stochastic delay most pronounced in the Finke–Watzky model?
These deviations become particularly pronounced when the “nucleation” step [Eq. (1)] is much slower (and thus much less likely) than the autocatalytic self-replicating reaction [Eq. (2)], and when no $B$ molecules are present initially., i.e.,
@@ III.1.3 Qualitative explanation of the observed effects
When the network (1)–(2) is initialized with $N_{B}(0)=0$, the waiting time for the first $\mathrm{A}\rightarrow\mathrm{B}$ event may be long. Because the waiting time to the first reaction is exponentially distributed (see Subsection III.1.5), many realizations remain in the initial state for a substantial period before any reaction occurs, whereas deterministic kinetics proceeds without delay. Consequently, the ensemble mean exhibits a systematic delay that can be captured only by solving the full CME or by stochastic simulation. The same reasoning can be applied to the case $N_{B}(0)=1$, although in this case it is much less compelling.
The origin of the large fluctuations around the mean for small $N_{B}(0)$ is likewise intuitive. Although the total molecule number is relatively large ($M=100$), the $B$-subsystem is initially very small (in the extreme case, absent). As a result, $N_{B}(t)$ shows large relative fluctuations; by the constraint (5), the same holds for $N_{A}(t)$.
While suggestive, these arguments are qualitative and incomplete, and may be misleading. A more careful analysis is therefore required.
Unfortunately, we are unable to provide a simple and general explanation for the discrepancies between deterministic and stochastic kinetics of the FWM (1)–(2). This situation is reminiscent of that in Ref. Schuster (2019), which likewise did not account for the origin of the “stochastic delay” in the case of the simple autocatalytic reaction (2). A partial understanding can, however, be obtained from the analysis of the moment equations (see the next subsection). In addition, Sec. III.1.5 clarifies the origin of the sensitivity of the CME solution to the initial conditions for selected parameter values.
@@
For simplicity, we now set $V=1$, so that $a_{s}(t)$ [Eq. (26)] reduces to $\mu_{1A}$, while the deterministic (macroscopic) rate constants $\mathcal{K}_{i}$, $i=1,2$, coincide with the mesoscopic rate constants $k_{i}$ of the CME. In this case, the deterministic concentration $a(t)$ evolves according to
@@
$\forall t\in[0,\infty):\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \dot{a}(t)\leq 0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \dot{\mu}_{1A}(t)\leq 0.$ (32)
@@
III.1.5 Sensitivity of stochastic dynamics to the initial number of B molecules
The stochastic delay resists a simple, complete, and general explanation. However, the pronounced dependence of the solution of the CME on the initial number of $B$ molecules, $N_{B}(0)$ (Figures 12), admits a more straightforward interpretation.
Specifically, for $N_{B}(0)=0$ ($N=M$), the transition rate $r_{N}$ from the initial state $n=N$ to n=N1n=N-1 is much smaller than in the case $N_{B}(0)=1$ (N=M1N=M-1). Consequently, the average waiting time for this first transition is significantly longer when $N_{B}(0)=0$. These intuitive observations can, however, be cast into a precise mathematical form, which we now proceed to develop as generally as possible.
Consider a set of irreversible chemical reactions involving species A and B, whose stochastic kinetics are described by the CME of the form (6) with the deterministic initial condition (7): we know with certainty that at $t=0$ the system contains $N_{A}(0)=N$ molecules of A and $N_{B}(0)=M-N$ molecules of B. The form of the CME (6) implies that the probability, starting from $n=N$, that no reaction has occurred by time $t$ (i.e., no transition from $n=N$ to n=N1n=N-1) is
$S_{N}(t|N,M)=e^{-r_{N}(M)t}=P_{N}(t|N,M),$ (44)
where $P_{N}(t|N,M)$ is given by (10). (In this subsection we explicitly denote the dependence of the relevant quantities on $N$ and $M$). Thus, $S_{N}(t|N,M)$ is the survival probability of state $n=N$ provided that the total number of molecules is $M$. Consequently, the probability density for the $N\to N-1$ transition occurring in the interval $[t,t+dt]$ is exponential:
$g(t_{1}|N,M)=-\frac{dS_{N}(t|N,M)}{dt}\Big{|}_{t=t_{1}}=r_{N}(M)e^{-r_{N}(M)t_{1}},$ (45)
where $t_{1}\in(0,t)$ is the time of the first reaction event. Its average value is equal
$\overline{t}_{1}(N,M)=\int_{0}^{\infty}t_{1}\,g(t_{1}|N,M)\,dt_{1}=\frac{1}{r_{N}(M)}.$ (46)
To proceed, we consider the probabilities $P_{n}(t|N,M)$, which are the solutions of the CME with the initial condition (7). For $n=1,2,\ldots,N-1$, these satisfy the identity
$P_{n}(t|N,M)=\int_{0}^{\infty}g(t_{1}|N,M)\,P_{n}(t-t_{1}|N-1,M)\,dt_{1},$ (47)
while $P_{N}(t|N,M)$ is given by (44). Equation (47) follows from the fact that the Laplace transform of a convolution is the product of Laplace transforms, and from the explicit form of the Laplace transform of $P_{n}(t|N,M)$, see Eqs. (83) and (84) in Appendix A.
Equation (47) implies the following expression for the average number of A molecules,
$\mu_{1A}(t|N,M)$ $=$ $\sum_{n=0}^{N}n\,P_{n}(t|N,M)=Ne^{-r_{N}(M)t}$
$+$ $\int_{0}^{\infty}g(t_{1}|N,M)\,\mu_{1A}(t-t_{1}|N-1,M)\,dt_{1}.$
Eq. (III.1.5) holds for any initial value $N$ of A molecules and total number $M$ of molecules.
Now, compare Eq. (III.1.5) in two cases: $N=M$ and N=M1N=M-1. If $r_{M}(M)\ll r_{M-1}(M)$, then $g(t_{1}|M,M)$ is much broader and flatter than $g(t_{1}|M-1,M)$. Therefore, the time evolution $\mu_{1A}(t|M,M)$, initially dominated by the term $Me^{-r_{M}(M)t}$, is much slower than the time evolution of $\mu_{1A}(t|M-1,M)$.
Let us now return to the Finke–Watzky model and to the situations illustrated in Figs. 1 and 2. For $N=M$, we have $r_{M}(M)=k_{1}M\ll r_{M-1}(M)=(M-1)(k_{1}+k_{2})$. For broad range of the model parameters, in particular for $M=100$ and $k_{1}=10^{-3}$, the distribution $g(t_{1}|M,M)$ given by Eq. (45) is very broad, and the term $Me^{-r_{N}(M)t}$ in Eq. (III.1.5) decays slowly. This explains why $\mu_{1A}(t|M,M)$ in Fig. 1 is significantly delayed relative to $\mu_{1A}(t|M-1,M)$ in Fig. 2. By choosing sufficiently small $k_{1}$, this delay can be made arbitrarily large. Note also that in this situation the average waiting time (46) is much larger for $N=M$ than for N=M1N=M-1:
$\overline{t}_{1}(M,M)=\frac{1}{k_{1}M}\gg\overline{t}_{1}(M-1,M)=\frac{1}{(M-1)(k_{1}+k_{2})}.$ (49)
As follows from Eq. (49), the same remains true when we compare $\overline{t}_{1}(M,M)$ for two different values of $k_{1}$: if $k^{\prime\prime}\ll k^{\prime}$, then $\overline{t}_{1}(M,M;k^{\prime\prime})\gg\overline{t}_{1}(M,M;k^{\prime})$. For $k^{\prime\prime}=10^{-4}$ and $k^{\prime}=10^{-3}$, this effect can be seen by comparing Fig. 1 and Fig. 6.
We emphasize that sensitivity of the CME solution to the initial condition is not specific to the FWM but can be expected in any network of irreversible reactions with the following properties: (i) it involves only two chemical species, A and B; (ii) the net result of each reaction in the network is the disappearance of one A molecule and the creation of one B molecule, so that the total number of molecules, $N_{A}(t)+N_{B}(t)=M$, is conserved; (iii) the transition rate for the first reaction event, $r_{N}$, is much smaller for $N=M$ than for N=M1N=M-1; (iv) the sequence $r_{n}$ decreases monotonically near $n=M$. Therefore, the above considerations also apply to self-replicating systems in which an improbable "spontaneous" process, modeled as a (pseudo) first-order reaction, is followed by autocatalytic replication. The latter need not be restricted to the form (2); it may involve higher-order autocatalysis, such as $\mathrm{A}+2\mathrm{B}\rightarrow 3\mathrm{B}$.
@@ III.2.1 Stochastic delay in the reaction network (3)–(4)
Figure 9: Stochastic versus deterministic kinetics for the reversible Finke–Watzky model (3)–(4) with $M=N=100$ ($N_{B}(0)=0$), $k_{1}^{+}=10^{-3}$, $k_{2}^{+}=k_{1}^{-}=10^{-1}$, and $k_{2}^{-}=2.0$. The deterministic concentration of A, $a(t)$ (105, red dashed), decreases faster than the stochastic mean $a_{s}(t)\equiv\mu_{1A}(t)/V$ (26, thick blue). Thin lines show $\mu_{1A}(t)\pm\sigma(t)$ as a measure of fluctuations. $V=1$ so that $a_{s}(t)=\mu_{1A}(t)$. All quantities are plotted versus $\tau=k_{2}t$. Even though, compared to Fig. 8, the equilibrium is shifted strongly toward the substrates, the stochastic delay remains large, as do the fluctuations around the mean number of A molecules. Despite the different dynamics of $a(t)$ and $\mu_{1A}(t)$, the stationary values coincide; this is not visible in the figure because the time axis is restricted to better highlight the differences in dynamics.
@@ III.2.2 Steady-state solution of the CME for the reactions (3)-(4)
$\tilde{M}=M+\frac{k^{+}_{1}}{k^{+}_{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{L}\equiv M+\frac{2k^{-}_{1}}{k^{-}_{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \theta\equiv\frac{k^{-}_{2}}{2k^{+}_{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \eta\equiv\frac{k^{-}_{1}}{k^{+}_{2}}$. (53)
@@ III.3 Two Reaction Networks Closely Related to the Finke–Watzky Model
In the case of the FWM [Eqs. (1)–(2)], for $N\geq\lfloor\tilde{M}/2\rfloor$, the sequence $\{r_{n}\}$ is non-monotonic: it increases for $n\leq\lfloor\tilde{M}/2\rfloor$ (or $n\leq\lfloor\tilde{M}/2\rfloor+1$) and then decreases, assuming $\tilde{M}\approx M$, which is the most relevant situation from our viewpoint. In particular, for $N=M$ the difference between $r_{N}$ and $r_{N-1}$ can be very substantial (see Sec. III.1.5).
@@
Figure 10: Stochastic versus deterministic kinetics for the reactions (64)–(65), with the initial condition $N_{B}(0)=0$ and rate constants $k_{1}=0.001$ and $k_{2}=0.1$. The deterministic concentration $a(t)$, given by Eq. (107), is plotted as a red dashed line. Its stochastic counterpart, $a_{s}(t)=\mu_{1A}(t)/V$ [Eq. (26)], is shown as a thick blue solid line. No visible difference is observed between the two time evolutions. Light blue dotted lines represent the single–standard deviation envelope around the mean, $\mu_{1A}(t)\pm\sigma(t)$, thus quantifying stochastic fluctuations, which are now much smaller than in the case of the FWM (1)–(2) for the same model parameters. The system volume is set to $V=1$, so that $k_{2}=\mathcal{K}_{2}$ and $a_{s}(t)=\mu_{1A}(t)$ with $\mu_{1A}(t)$ given by Eq. (III.1.1). All quantities are plotted as functions of the dimensionless time variable $\tau=k_{2}t$.

### v3

@@ I Introduction
The stochastic delay and the fluctuations around the mean are especially significant when the system is initially free of $\mathrm{B}$ and when $k_{1}\ll k_{2}$. For this reason, unlike in Ref. Schuster (2019), we do not use SSA to compute the stochastic dynamics of the system—Gillespie’s standard algorithm becomes ineffective when reaction rates differ greatly, and advanced variants such as tau-leaping Gillespie (2001); Cao et al. (2005, 2006, 2007) are not fully equivalent to the CME. Therefore, we derive an explicit analytical solution of the CME for the reactions (1)–(2). We are not aware of any prior publication of such a result.
@@ II Theoretical Framework
We assume the system under consideration to be spatially homogeneous and closed, with no exchange of molecules with the environment (e.g., well‑stirred reagents in a batch reactor). Consequently, the total molecule number is conserved:
@@ II.1 Irreversible reactions
Both the Finke–Watzky model (1)–(2) and the two other sets of irreversible reactions considered here are one‑step processes. Their Chemical Master Equation (CME) takes the general form van Kampen (2007):
@@
for $n=1,\dots,N-1$. $P_{0}(t)$ follows from normalization or directly from Eq. (6) at $n=0$. Here $C_{NN}=1$. Finally, since $r_{0}=0$ and $r_{n}>0$ for $n\geq 1$, the steady‐state solution is
@@
The general analytical solution (10)–(11) of Eqs. (6) has appeared in various contexts and has been derived independently by different authors Allnatt and Jacobs (1968); Arslan and Laurenzi (2008); Lee and Kim (2012); Rudnicki (2014); Jedrak (2014). In contrast, by "explicit analytical solution" we mean here not the general form (10)–(11) of the solution, but the most concise analytical form of both $P_{n}(t)$ and its lowest moments that can be obtained for a specific choice of the coefficients $r_{n}$.
@@ II.2 Reversible reactions
For reversible bimolecular reactions in general, and for (3)–(4) in particular, the functional form of $\mathbf{P}(t)$ can no longer be determined analytically, and we must resort to numerical methods. This may involve either SSA simulations or a direct, exact numerical solution of the system of ODEs (14) or a method of Ref. Smith and Shahrezaei (2015).
@@ III Results
In this section, we present our results. First, we derive the explicit analytical solution of the CME for the original formulation of the Finke–Watzky model (FWM) (1)–(2). We then use this solution to analyze the differences between stochastic and deterministic kinetics for selected choices of model parameters and initial conditions (Subsection III.1.2). The approximate approach based on a truncated set of moment equations is discussed in Subsection III.1.4.
@@
$\tilde{M}\equiv M+\nu,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \nu\equiv\frac{k_{1}}{k_{2}},$ (20)
@@ III.1.1 Probability distribution and its two lowest moments
Analytical expressions for the moments (III.1.1)–(III.1.1) reduce the double sum (which arises when calculating the moment directly from its definition using $P_{n}(t)$ given by Eq. 22) to a single sum, lowering the computational cost. A caveat is that numerical evaluation of Eqs. (III.1.1) and (III.1.1) requires arbitrary-precision arithmetic: for $M=100$ the absolute values of the coefficients $E^{(1)}_{l}$ and $E^{(2)}_{l}$ reach $10^{30}$ and $10^{33}$, respectively, so small relative round-off errors at short times can cause large absolute errors in the moments. [^3] Here, the GNU Multiple Precision Floating-Point Reliable Library (GNU MPFR) was used for this purpose (https://www.mpfr.org). In fact, instead of relying on the potentially cumbersome arbitrary-precision evaluation of Eqs. (III.1.1)–(III.1.1), one can simply solve the set of $N+1$ ODEs in Eq. (6) by direct numerical integration. We performed the calculations using both approaches and obtained perfect agreement with the analytical solution.
@@ III.1.2 When is the stochastic delay most pronounced in the Finke–Watzky model?
These deviations become particularly pronounced when the “nucleation” step [Eq. (1)] is much slower (and thus much less likely) than the autocatalytic reaction [Eq. (2)], and when no $B$ molecules are present initially., i.e.,
@@ III.1.3 Qualitative explanation of the observed effects
When the network (1)–(2) is initialized with $N_{B}(0)=0$, the waiting time for the first $\mathrm{A}\rightarrow\mathrm{B}$ event may be long. Because this waiting time is exponentially distributed, many realizations remain in the initial state for a substantial period before any reaction occurs, whereas deterministic kinetics proceeds without such a delay. Consequently, the ensemble mean exhibits a systematic lag that can be captured only by solving the full CME or by stochastic simulation. The same reasoning can be applied to the case $N_{B}(0)=1$, although here it is much less compelling.
The explanation of the strong dependence of the CME solution on the initial number of $B$ molecules, $N_{B}(0)$ (Figures 12), is similar. For $N_{B}(0)=0$ ($N=M$), the transition rate $r_{N}$ from the initial state $n=N$ to n=N1n=N-1 is much smaller than in the case $N_{B}(0)=1$ (N=M1N=M-1). Consequently, the average waiting time for this first transition is significantly longer when $N_{B}(0)=0$.
The origin of the large fluctuations around the mean for small $N_{B}(0)$ is likewise intuitive. Although the total molecule number is relatively large ($M=100$), the $B$-subsystem is initially very small (in the extreme case, absent). As a result, $N_{B}(t)$ shows large relative fluctuations; by the constraint (5), the same holds for $N_{A}(t)$. While suggestive, these arguments are qualitative and incomplete, and may be misleading. A more careful analysis is therefore required.
A full explanation of the discrepancies between deterministic and stochastic kinetics in the FWM (1)–(2) and related autocatalytic networks is beyond the scope of this work. Here, we focus instead on illustrating and comparing the exact stochastic and deterministic solutions for this particular reaction network, for which an explicit analytical solution of the CME can be obtained. A partial understanding of the origin of the stochastic delay can, however, be obtained from an analysis of the moment equations (see the next subsection).
@@
For simplicity, we now set $V=1$, so that $a_{s}(t)$ [Eq. (26)] reduces to $\mu_{1A}$, $a_{0}=N$, $b_{0}=M-N$, while the deterministic (macroscopic) rate constants $\mathcal{K}_{i}$, $i=1,2$, coincide with the mesoscopic rate constants $k_{i}$ of the CME. In this case, the deterministic concentration $a(t)$ evolves according to
@@
$\forall t\in[0,\infty):\penalty 10000\ \penalty 10000\ \penalty 10000\ \dot{a}(t)\leq 0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \dot{\mu}_{1A}(t)\leq 0.$ (32)
@@ III.2.1 Stochastic delay in the reaction network (3)–(4)
Figure 9: Stochastic versus deterministic kinetics for the reversible Finke–Watzky model (3)–(4) with $M=N=100$ ($N_{B}(0)=0$), $k_{1}^{+}=10^{-3}$, $k_{2}^{+}=k_{1}^{-}=10^{-1}$, and $k_{2}^{-}=2.0$. The deterministic concentration of A, $a(t)$ (99, red dashed), decreases faster than the stochastic mean $a_{s}(t)\equiv\mu_{1A}(t)/V$ (26, thick blue). Thin lines show $\mu_{1A}(t)\pm\sigma(t)$ as a measure of fluctuations. $V=1$ so that $a_{s}(t)=\mu_{1A}(t)$. All quantities are plotted versus $\tau=k_{2}t$. Even though, compared to Fig. 8, the equilibrium is shifted strongly toward the substrates, the stochastic delay remains large, as do the fluctuations around the mean number of A molecules. Despite the different dynamics of $a(t)$ and $\mu_{1A}(t)$, the stationary values coincide; this is not visible in the figure because the time axis is restricted to better highlight the differences in dynamics.
@@ III.2.2 Steady-state solution of the CME for the reactions (3)-(4)
$\tilde{M}=M+\frac{k^{+}_{1}}{k^{+}_{2}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \tilde{L}\equiv M+\frac{2k^{-}_{1}}{k^{-}_{2}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \theta\equiv\frac{k^{-}_{2}}{2k^{+}_{2}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \eta\equiv\frac{k^{-}_{1}}{k^{+}_{2}}$. (47)
@@ III.3 Two Reaction Networks Closely Related to the Finke–Watzky Model
In the case of the FWM [Eqs. (1)–(2)], for $N\geq\lfloor\tilde{M}/2\rfloor$, the sequence $\{r_{n}\}$ is non-monotonic: it increases for $n\leq\lfloor\tilde{M}/2\rfloor$ (or $n\leq\lfloor\tilde{M}/2\rfloor+1$) and then decreases, assuming $\tilde{M}\approx M$, which is the most relevant situation from our viewpoint. In particular, for $N=M$ the difference between $r_{N}$ and $r_{N-1}$ can be very substantial.
@@
Figure 10: Stochastic versus deterministic kinetics for the reactions (58)–(59), with the initial condition $N_{B}(0)=0$ and rate constants $k_{1}=0.001$ and $k_{2}=0.1$. The deterministic concentration $a(t)$, given by Eq. (101), is plotted as a red dashed line. Its stochastic counterpart, $a_{s}(t)=\mu_{1A}(t)/V$ [Eq. (26)], is shown as a thick blue solid line. No visible difference is observed between the two time evolutions. Light blue dotted lines represent the single–standard deviation envelope around the mean, $\mu_{1A}(t)\pm\sigma(t)$, thus quantifying stochastic fluctuations, which are now much smaller than in the case of the FWM (1)–(2) for the same model parameters. The system volume is set to $V=1$, so that $k_{2}=2\mathcal{K}_{2}$ (see Eq. (102)) and $a_{s}(t)=\mu_{1A}(t)$ with $\mu_{1A}(t)$ given by Eq. (III.1.1). All quantities are plotted as functions of the dimensionless time variable $\tau=k_{2}t$.